OpenStudy (anonymous):

Find (2n-3)(n-4) please help me, i am not that good at math and i dont have much time left to complete :(

OpenStudy (anonymous):

 or\[\left( 2 n-3 \right)\left( n-4 \right)=2 n^2-8 n-3 n+12=2 n^2-11 n+12\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw